# Dual axis atomic magnetometer and gyroscope enabled by nuclear spin perturbation

### v2

@@ Abstract
Atom beam experiment; qubit manipulation; Rb atom beam
@@ I Introduction
To distinguish the target signal from external magnetic noise, a common method is the Nuclear Magnetic Resonance (NMR) gyroscope which uses two noble gas species with different nuclear gyromagnetic ratios [13]. The two species make nominally independent measurements of the effective magnetic fields they see, and a discrepancy in the measurements reveals the rotation rate. The role of the alkali atoms in this scheme is to pump the noble gas species and measure their dynamics, via the magnetic interaction between $\mathbf{P^{n}}$ and $\mathbf{P^{e}}$. Recent work has also shown that modulation of the pump and applied bias field in NMR systems may be useful for reducing technical noise [14, 15]. The alkali in an NMR gyroscope does not work in the SERF regime, however, so spin relaxation noise limits how well an alkali species can measure the precession of the noble gas. To partly mitigate spin relaxation, pump pulses with alternating polarisation [16] and magnetic pulses [17] have shown promise.
@@
In this work, we introduce a Perturbed Alkali-Noble Comagnetometer (PANCo) scheme. It may may be considered an evolution of a SCC, although self-compensation is no longer required.
@@ II Methods
Before diving into a model of our scheme, it is helpful to have a picture of how it works. Without application of magnetic pulses, consider a constant pure rotation in the $x$-axis. Both the alkali and noble gas species rotate together. The optical pump will reset $\mathbf{P^{e}}$ to the $z$-axis and it will then precess around $\mathbf{P^{n}}$, which will remain displaced from the $z$-axis. The direct effect of the rotation on $\mathbf{P^{e}}$ may be neglected as the large gyromagnetic ratio of the alkali ensures that its magnetic response to $\mathbf{P^{n}}$ is dominant. Now consider the case of a pure magnetic field along the $x$-axis. This will also result in a movement of $\mathbf{P^{n}}$. When the alkali is pumped back to the $z$-axis, however, $\mathbf{P^{e}}$ will be now be driven by the combination of $\mathbf{P^{n}}$ and the external magnetic field. Despite the different origins, rotation and magnetic driving both lead to precession of $\mathbf{P^{e}}$ around some effective magnetic field and there is no way to discern the origin of that effective field.
@@ II.1 Model
To model our method we can use the coupled Bloch equations for the spin polarisation of the alkali atoms, $\mathbf{P^{e}}$, and the noble gas atoms, $\mathbf{P^{n}}$. The superscripts reflect the fact that the alkali atom spin is dominated by the electron spin while the noble gas polarisation is due to it’s nuclear spin. The equations of motion can be written as [12]
@@
where $\mathbf{B}$ is an external magnetic field, $\bm{\Omega}$ the angular velocity of rotation, $\gamma_{e}$ is the electron gyromagnetic ratio of the alkali metal, $\gamma_{n}$ is the nuclear gyromagnetic ratio of noble gas. Adding to the effect of the magnetic field there is also the magnetisation of the atomic species. The motion of $\mathbf{P^{e}}$ and $\mathbf{P^{n}}$ are driven by $\lambda M^{n}\mathbf{P^{n}}$ and $\lambda M^{e}\mathbf{P^{n}}$ respectively where $\lambda$ is the coupling enhancement between the polarisations, and $M^{(n,e)}$ are the maximum magnetisations of the atomic species. Terms that lead to decay of the polarisations are governed by $R_{se}^{(en,ne)}$ and $R_{c}^{(n,e)}$. These describe the rates of polarisation transfer from alkali to noble gas and noble gas to alkali as well as the spin-destroying collisions experienced by the noble gas and alkali species. Growth in the polarisation is due to the pumping rate, $R_{p}$, which is due to the pump beam that propagates in the direction of the pumping vector, $\mathbf{s}$. In the SERF regime there is a slowing down factor, $q$ that results from the nuclear angular momentum [8]. The slowing-down factor is a function of the polarisation of the alkali atoms. As our system is pulsed, $q$ will be varying throughout the readout phase as the alkali polarisation decays.
@@
The data in Fig. 2(a-d) show the evolution of both $\mathbf{P^{e}}$ and $\mathbf{P^{n}}$ projected onto the $x-y$ plane. In each case, $\mathbf{P^{n}}$ alternates between two points related by a $\pi/2$ rotation about the $z$-axis (which points out of the page), as it reacts to the alternating magnetic pulses. During the measurement of the alkali precession $\mathbf{P^{n}}$ is nearly constant. The alkali atoms, on the other hand, begin each decay period $\mathbf{P^{e}}$ aligned with the $z$-axis due to the optical pumping, and then undergo free precession.
@@
In an experiment we project only one axis of $\mathbf{P^{e}}$ into the optical polarisation of the probe beam. In this example we have measured the $x$-axis polarisation and plotted it in part Fig. 2(e). Looking at only the first pulse response, up to 10 ms, the traces look quite similar, but looking at the combined response over two alternate magnetic pulses we see responses for each of the rotations and magnetic fields that are substantially different. For the small signals normally measured with these devices, our system is linear, so that any combination of $x$ and $y$ rotations and magnetic fields can be found using linear regression over the four ‘signature’ traces shown in Fig. 2(e). By performing this fit for each measurement cycle, our method can be used to extract time series for the four signals. This processing is not intensive and can be done in real-time with minimal computational resources.
@@ III Experimental realisation
We have conducted a proof of principle experiment to demonstrate the feasibility of our scheme. The cell used was approximately spherical and 1cm in diameter. It’s nominal room-temperature contents where 0.01 atm of ${}^{129}\mathrm{Xe}$, 0.4 atm of N2, and a drop of ${}^{87}\mathrm{Rb}$. This was heated to 125 degrees Celsius inside a four-layer mu-metal shield (Twinleaf MS1) with built in 3-axis magnetic coils. The pump light was provided by a tapered amplifier seeded by a 795 nm distributed Bragg reflector laser (Photodigm). The current to the tapered amplifier was pulsed to form the pump pulses. This provided around 2 W peak power while on, with a leakage of the seed laser of approximately 10 $\mu$W with the current off. The beam was enlarged to be similar to the cell size at roughly 10 mm $\times$ 10 mm. The pump pulses were 300 $\mu$s and applied every 2 ms. After several minutes of equilibrating time, there was a $\sim$100 nT effective field on the alkali due to ${}^{129}\mathrm{Xe}$ polarisation. The probe light was provided by a 780 nm external cavity diode laser, detuned from the D2 line by $\sim$100 GHz and with a power of 2 mW. In these conditions, the T2 times of ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ were measured to be near 2 ms and 3 s respectively.
@@
To set the operating point, a simple procedure was followed. The bias field was tuned to give approximately half of a Larmor period within the 1.7 ms free decay window. This resulted in a bias field significantly below the compensation point. The transverse fields were adjusted to minimise the size of the residual signal. The PANCo magnetic pulses were then turned on. To set the pulse amplitude, a 100 pT transverse field was applied while the pulse amplitude was increased from zero until the difference in signal between alternate windows was similar in size to the signal in each window.
Figure 3: The polarisation of the ${}^{87}\mathrm{Rb}$ ensemble measured in response to magnetic fields perpendicular to the pump axis. As in Fig. 2(e), this data is measured over one cycle of the alternating magnetic pulses. a) Response to steady magnetic fields. These curves provide us with to the magnetic field responses $S^{B}_{x,y}$, similar to those modelled in Fig. 2. b) Response immediately following application of a magnetic field step. As described in the text, these curves allow the extraction of a proxy for non-magnetic (rotational) signals in our data. These signatures were measured using $\approx 100\,pT$ amplitude square wave driving along the x- and y-axes of the system.
@@
The signatures for $B_{x,y}$ are obtained easily since the coils in the oven allow precise application of magnetic fields. The measured curves for $S^{B}_{x,y}$ are shown in Fig. 3(a). As we are lacking a precision rotation stage, however, we are unable to easily extract $S^{\Omega}_{x,y}$. Instead, we use the response to a magnetic square wave to extract a proxy for rotational signals.
We can consider the response of $\mathbf{P^{e}}$ to a magnetic field to be the sum of two components, $S^{B}_{x,y}=S^{B,direct}_{x,y}+S^{B,n}_{x,y}$, where the two terms on the right are the alkali’s direct response to the applied field and that due to the field of the noble gas, respectively. We may obtain a good measurement of $S^{B,direct}_{x,y}$ by measuring $\mathbf{P^{e}}$ following a fast magnetic step, before the noble gas can respond. Now we recall that the system response to a rotation is dominated by the movement of $\mathbf{P^{n}}$. Since $\mathbf{P^{e}}$ does not care what has caused displacement of $\mathbf{P^{n}}$, there must be a relationship between $S^{B,n}_{x,y}$ and $S^{\Omega}_{x,y}$. In fact, in the regime we consider here, they are simply scaled versions of each other, as can be seen from Fig. 2(a,b) and the associated discussion. Thus, as a proxy for rotation or other non-magnetic signals we use a signature of $S^{B,n}_{x,y}=S^{B}_{x,y}-S^{B,direct}_{x,y}$.
Based on the above considerations, we can now make use of the data from square wave modulation to extract our rotational proxy. At the end of the square wave, once the magnetic field has been stable for sufficient time, the experiment will be measuring $S^{B}_{x,y}$, the steady state magnetic signature, which is shown in Fig. 3(a). Immediately after the step change in magnetic field, on the other hand, the alkali will respond only to the applied magnetic field, because the nuclear spin had not yet precessed due to its slow response. The precession of $\mathbf{P^{e}}$ after the step change therefore gives us $S^{B,direct}_{x,y}$, which is shown in Fig. 3(b). We can then combine these data to find $S^{B,n}_{x,y}=S^{B}_{x,y}-S^{B,direct}_{x,y}$, which as noted above, is proportional to the rotational signature $S^{\Omega}_{x,y}$. We will refer to signals extracted using this signature as $\beta(x,y)$.
@@
The results of using our measured and inferred signatures are shown in Fig. 4. Square wave magnetic modulations were applied on each axis. Every 4 ms, pairs of precession periods were recorded and then fitted using the signatures. In Fig. 4(a) we fitted using $S^{B}_{x,y}$ to find $B_{x}$ and $B_{y}$, giving the expected square wave modulation. In Fig. 4(b) we fitted the precession periods using $S^{B,n}_{x,y}$ to find a rotational or non-magnetic, signal that we call $\beta$. This shows spikes in response to the rapid change in magnetic field before the noble gas has time to respond. This is qualitatively the same behaviour seen in CW comagnetometers.
@@ IV Idealised Sensitivity
We now consider the limits of the PANCo scheme independent of experimental issues, as an estimate of how well the scheme might perform when compared to state-of-the-art experiments. We also consider optimising performance in terms of bias field.
@@
Figure 7: The sensitivity of PANCo, normalised to the optimum performance of CW systems. (a) Comparison of the PANCo scheme rotational sensitivity as a function of bias field under white noise conditions, normalised to the optimal performance of a single axis CW SCC with the same parameters. The PANCo reaches optimum sensitivity at bias of 106.3 nT, marked with a dashed line. For reference, the sensitivity of the CW SCC is also shown, although it can only suppress magnetic fields at the compensation point of 104.5 nT, marked with a dashed line. (b) The PANCo scheme measures transverse magnetic fields simultaneously with rotation. The magnetic sensitivity has been scaled to the performance of a SERF alkali magnetometer that was modelled by removing the noble gas and setting the bias field to zero. At 106.3 nT, the magnetic sensitivity of PANCo is a factor of four worse than the ideal alkali magnetometer. The peaks near 100 nT are due to the alkali seeing near-zero field such that the 4 signatures are no longer distinguishable.
@@ V Discussion
Working towards these applications will require careful consideration of the magnetic cross-talk. The PANCo scheme can theoretically achieve zero cross-talk between magnetic and rotation signals, although this will only hold true when the signatures are known perfectly. Achieving low cross-talk in practice means good measurement of the signatures in the first place, and then low drift of experimental parameters so the signatures do not change. A common issue in comagnetometers is drift of the compensation point due either to stray $B_{z}$ fields or drifting polarisation. For the simulation at 106.3 nT we calculate that the bias field must remain within ±$\pm 0.2\,$nT to keep the cross-talk below $0.2\,\mu$Hz/pT. This value of cross-talk corresponds to a ‘suppression’ factor of around 150 for an equivalent CW comagnetometer. This corresponds to a relative bias stability of 0.2% which is stringent, but achievable. Further, drifts in $B_{z}$ are easily detected in our scheme via the alkali precession frequency (by introducing a small transverse field if necessary), thus enabling a method for stabilising the bias field. We also expect drifts in other parameters, such as the magnetization of the spins, although the use of pulsed pumping will help in distinguishing this and other deleterious effects. It is notable that our experiment, albeit with a different mixture of gases, has already achieved a high level of magnetic suppression with only minimal optimisation. This gives us reason to be optimistic about the levels of suppression that can be achieved in future work.
@@
The large magnetic pulses introduced in the PANCo scheme are also a significant potential noise source. We expect that variations in pulse strength will cause variations in the signatures, and thus cross-talk between magnetic and rotation signals similar to variation in bias field. In addition, these large pulses could induce eddy currents in nearby conductors such as magnetic shields, which may in turn cause magnetic noise that persists during the free precession period. Variations in these eddy currents could be caused by pulse size drifts or external effects such as temperature changes, and so cause real and unwanted variations in magnetic fields. At present neither of these effects seem dominant in our experiment: we conclude this from a simple test with 10$\times$ weaker pulses (area of $\pi/20$, obtained by turning down the amplifier). This gave no corresponding reduction in low frequency magnetic noise floor, and only a degradation in rotational noise floor due to degraded magnetic suppression. However we expect the tuning and control of the magnetic pulses to be a consideration for future work. It is perhaps salient to note that the application of phase-controlling pulses are used in quantum information processing and NMR for dynamic decoupling [33, 34, 35]. While the purpose of the pulses in our system is different, there is a wealth of experimental and theoretical work that shows pulse sequences can be used in quantum limited systems with great success.

### v3

@@ I Introduction
To distinguish the target signal from external magnetic noise, a common method is the Nuclear Magnetic Resonance (NMR) gyroscope which uses two noble gas species with different nuclear gyromagnetic ratios [31]. The two species make nominally independent measurements of the effective magnetic fields they see, and a discrepancy in the measurements reveals the rotation rate. The role of the alkali atoms in this scheme is to pump the noble gas species and measure their dynamics, via the magnetic interaction between $\mathbf{P^{n}}$ and $\mathbf{P^{e}}$. Recent work has also shown that modulation of the pump and applied bias field in NMR systems may be useful for reducing technical noise [30, 27]. The alkali in an NMR gyroscope does not work in the SERF regime, however, meaning that spin relaxation noise limits how well an alkali species can measure the precession of the noble gas. To partly mitigate spin relaxation, pump pulses with alternating polarisation [18] and magnetic pulses [14] have shown promise.
@@
In this work, we introduce a Perturbed Alkali-Noble Comagnetometer (PANCo) scheme. It may be considered an evolution of a SCC, although self-compensation is no longer required.
@@ II Methods
Before diving into a model of our scheme, it is helpful to have a picture of how it works. Without application of magnetic pulses, consider a constant pure rotation in the $x$-axis. Both the alkali and noble gas species rotate together. The optical pump will reset $\mathbf{P^{e}}$ to the $z$-axis and it will then precess around $\mathbf{P^{n}}$, which will remain displaced from the $z$-axis. The direct effect of the rotation on $\mathbf{P^{e}}$ may be neglected as the large gyromagnetic ratio of the alkali ensures that its magnetic response to $\mathbf{P^{n}}$ is dominant. Now consider the case of a pure magnetic field along the $x$-axis. This will also result in a movement of $\mathbf{P^{n}}$. When the alkali is pumped back to the $z$-axis, however, $\mathbf{P^{e}}$ will now be driven by the combination of $\mathbf{P^{n}}$ and the external magnetic field. Despite the different origins, rotation and magnetic driving both lead to precession of $\mathbf{P^{e}}$ around some effective magnetic field and there is no way to discern the origin of that effective field.
@@ II.1 Model
To model our method we can use the coupled Bloch equations for the spin polarisation of the alkali atoms, $\mathbf{P^{e}}$, and the noble gas atoms, $\mathbf{P^{n}}$. The superscripts reflect the fact that the alkali atom spin is dominated by the electron spin while the noble gas polarisation is due to its nuclear spin. The equations of motion can be written as [13]
@@
where $\mathbf{B}$ is an external magnetic field, $\bm{\Omega}$ the angular velocity of rotation, $\gamma_{e}$ is the electron gyromagnetic ratio of the alkali metal, $\gamma_{n}$ is the nuclear gyromagnetic ratio of noble gas. Adding to the effect of the magnetic field there is also the magnetisation of the atomic species. The motion of $\mathbf{P^{e}}$ and $\mathbf{P^{n}}$ are driven by $\lambda M^{n}\mathbf{P^{n}}$ and $\lambda M^{e}\mathbf{P^{e}}$ respectively where $\lambda$ is the coupling enhancement between the polarisations, and $M^{(n,e)}$ are the maximum magnetisations of the atomic species. Terms that lead to decay of the polarisations are governed by $R_{se}^{(en,ne)}$ and $R_{c}^{(n,e)}$. These describe the rates of polarisation transfer from alkali to noble gas and noble gas to alkali as well as the spin-destroying collisions experienced by the noble gas and alkali species. Growth in the polarisation is given by the pumping rate, $R_{p}$, which results from the pump beam propagating along the $z$-axis. In the SERF regime there is a slowing down factor, $q$ that results from the nuclear angular momentum [1]. The slowing-down factor is a function of the polarisation of the alkali atoms. As our system is pulsed, $q$ will be varying throughout the readout phase as the alkali polarisation decays.
@@
The data in Fig. 2(a-d) show the evolution of both $\mathbf{P^{e}}$ and $\mathbf{P^{n}}$ projected onto the $x-y$ plane. In each case, $\mathbf{P^{n}}$ alternates between two points related by a $\pi/2$ rotation about the $z$-axis (which points out of the page), as it reacts to the alternating magnetic pulses. During the measurement of the alkali precession, $\mathbf{P^{n}}$ remains nearly constant. The alkali atoms, on the other hand, begin each decay period with $\mathbf{P^{e}}$ aligned to the $z$-axis due to the optical pumping, and then undergo free precession.
@@
In an experiment we project only one axis of $\mathbf{P^{e}}$ into the optical polarisation of the probe beam. In this example we have measured the $x$-axis polarisation and plotted it in Fig. 2(e). Looking at only the first pulse response, up to 10 ms, the traces look quite similar, but looking at the combined response over two alternate magnetic pulses we see responses for each of the rotations and magnetic fields that are substantially different. For the small signals normally measured with these devices, our system is linear, so that any combination of $x$ and $y$ rotations and magnetic fields can be found using linear regression over the four ‘signature’ traces shown in Fig. 2(e). By performing this fit for each measurement cycle, our method can be used to extract time series for the four signals. This processing is not intensive and can be done in real-time with minimal computational resources.
@@ III Experimental realisation
We have conducted a proof of principle experiment to demonstrate the feasibility of our scheme. The cell used was approximately spherical and 1cm in diameter. Its nominal room-temperature contents where 0.01 atm of ${}^{129}\mathrm{Xe}$, 0.4 atm of N2, and a drop of ${}^{87}\mathrm{Rb}$. This was heated to 125 degrees Celsius inside a four-layer mu-metal shield (Twinleaf MS1) with built in 3-axis magnetic coils. The pump light was provided by a tapered amplifier seeded by a 795 nm distributed Bragg reflector laser (Photodigm). The current to the tapered amplifier was pulsed to form the pump pulses. This provided around 2 W peak power while on, with a leakage of the seed laser of approximately 10 $\mu$W with the current off. The beam was enlarged to be similar to the cell size at roughly 10 mm $\times$ 10 mm. The pump pulses were 300 $\mu$s and applied every 2 ms. After several minutes of equilibrating time, there was a $\sim$100 nT effective field on the alkali due to ${}^{129}\mathrm{Xe}$ polarisation. The probe light was provided by a 780 nm external cavity diode laser, detuned from the D2 line by $\sim$100 GHz and with a power of 2 mW. In these conditions, the T2 times of ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ were measured to be near 2 ms and 3 s respectively.
@@
To set the operating point, a simple procedure was followed. The bias field was tuned to give approximately half of a Larmor period within the 1.7 ms free decay window. This resulted in a bias field significantly below the compensation point. At this point, a signal was observed due primarily to residual transverse fields and beam misalignments. This was minimized by applying small field corrections and beam pointing corrections. The PANCo magnetic pulses were then turned on. To set the pulse amplitude, a 100 pT transverse field was applied while the pulse amplitude was increased from zero until the difference in signal between alternate windows was similar in size to the signal in each window.
Figure 3: The polarisation of the ${}^{87}\mathrm{Rb}$ ensemble measured in response to magnetic fields perpendicular to the pump axis. As in Fig. 2(e), this data is measured over one cycle of the alternating magnetic pulses. a) Response to steady magnetic fields. These curves provide us with the magnetic field responses $S^{B}_{x,y}$, similar to those modelled in Fig. 2. b) Response immediately following application of a magnetic field step. As described in the text, these curves allow the extraction of a proxy for non-magnetic (rotational) signals in our data. These signatures were measured using $\approx 100\,pT$ amplitude square wave driving along the x- and y-axes of the system.
@@
The signatures for $B_{x,y}$ are obtained easily since the coils in the oven allow precise application of magnetic fields. The measured curves for $S^{B}_{x,y}$ are shown in Fig. 3(a). As we are lacking a precision rotation stage it is not possible to similarly extract $S^{\Omega}_{x,y}$. Instead, we use the response to a magnetic square wave to extract a proxy for rotational signals.
We can consider the response of $\mathbf{P^{e}}$ to a magnetic field to be the sum of two components, $S^{B}_{x,y}=S^{B,direct}_{x,y}+S^{B,n}_{x,y}$, where the two terms on the right are the alkali’s direct response to the applied field and that due to the field of the noble gas, respectively. We may first find $S^{B,direct}_{x,y}$ by measuring $\mathbf{P^{e}}$ immediately following a fast magnetic step, before the noble gas can respond. Next, we recall that the system response to a rotation is dominated by the movement of $\mathbf{P^{n}}$. Since $\mathbf{P^{e}}$ does not care what has caused displacement of $\mathbf{P^{n}}$, there must be a relationship between $S^{B,n}_{x,y}$ and $S^{\Omega}_{x,y}$. In fact, in the regime we consider here, they are simply scaled versions of each other, as can be seen from Fig. 2(a,b) and the associated discussion. Thus, as a proxy for rotation or other non-magnetic signals we use the signatures given by $S^{B,n}_{x,y}=S^{B}_{x,y}-S^{B,direct}_{x,y}$. These proxy signatures are not calibrated, although we note that a more complete analysis of a magnetic step change has been used to extract a calibrated rotational response of a CW CSS [22].
Based on the above considerations, we can now make use of the data from square wave modulation to extract our rotational proxies. At the end of the square wave, once the magnetic field has been stable for sufficient time, the experiment will be measuring $S^{B}_{x,y}$, the steady state magnetic signatures, which are shown in Fig. 3(a). Immediately after the step change in magnetic field, on the other hand, the alkali will respond only to the applied magnetic field, because the nuclear spin had not yet precessed due to its slow response. The precession of $\mathbf{P^{e}}$ after the step change therefore gives us $S^{B,direct}_{x,y}$, which are shown in Fig. 3(b). We can then combine these data to find $S^{B,n}_{x,y}=S^{B}_{x,y}-S^{B,direct}_{x,y}$, which as noted above, are proportional to the rotational signatures $S^{\Omega}_{x,y}$. We will refer to signals extracted using this signatures as $\beta_{(x,y)}$.
@@
The results of using our measured and inferred signatures are shown in Fig. 4. Square wave magnetic modulations were applied on each axis. Every 4 ms, pairs of precession periods were recorded and then fitted using the signatures. In Fig. 4(a) we show the fitted amplitudes of $B_{x}$ and $B_{y}$, giving the expected square wave modulation. In Fig. 4(b) we show the fitted amplitudes $\beta$ that correspond to rotational or non-magnetic signals. These curves show spikes in response to the rapid change in magnetic field before the noble gas has time to respond. This is qualitatively the same behaviour seen in CW comagnetometers.
@@ IV Idealised Sensitivity
We now consider the limits of the PANCo scheme independent of experimental issues, as an estimate of how well the scheme might perform when applied to state-of-the-art systems that typically use K-3He. We also consider optimising performance in terms of bias field.
@@
Figure 7: PANCo Sensitivity to rotations and magnetic fields. (a) Comparison of the PANCo rotational sensitivity as a function of bias field under white noise conditions. The sensitivity is compared to a single axis CW SCC with the same operating parameters. The sensitivities have been normalised to the best performance of the CW SCC, which occurs at the self-compensation point where the bias field is 104.5 nT, as marked by a dotted line. The PANCo achieves a sensitivity that is 30% better than the CW SCC at bias of 106.3 nT, also marked with a dotted line. We note that the model does not consider magnetic noise, which is only suppressed by the CW SCC at the self-compensation point. (b) The PANCo scheme measures transverse magnetic fields simultaneously with rotation. The magnetic sensitivity has been normalised to the performance of a SERF alkali magnetometer that was modelled by removing the noble gas and setting the bias field to zero. At 106.3 nT, the magnetic sensitivity of PANCo is a factor of four worse than the ideal alkali magnetometer. The peaks near 100 nT are due to the alkali seeing near-zero field such that the 4 signatures are no longer distinguishable.
@@ V Discussion
Working towards these applications will require careful consideration of the magnetic cross-talk. The PANCo scheme can theoretically achieve zero cross-talk between magnetic and rotation signals, although this will only hold true when the signatures are known perfectly. Achieving low cross-talk in practice means good measurement of the signatures in the first place, and then low drift of experimental parameters so the signatures do not change. A common issue in comagnetometers is drift of the compensation point due either to stray $B_{z}$ fields or drifting polarisation. For the simulation at 106.3 nT we calculate that the bias field must remain within ±$\pm 0.2\,$nT to keep the cross-talk below $0.2\,\mu$Hz/pT. This value of cross-talk corresponds to a ‘suppression’ factor of around 150 for an equivalent CW comagnetometer. This corresponds to a relative bias stability of 0.2% which is stringent, but achievable. Further, drifts in $B_{z}$ are easily detected in our scheme via the alkali precession frequency (by introducing a small transverse field if necessary), thus enabling a method for stabilising the bias field [10]. We also expect drifts in other parameters, such as the magnetization of the spins, although the use of pulsed pumping will help in distinguishing this and other deleterious effects. It is notable that our experiment, albeit with a different mixture of gases, has already achieved a high level of magnetic suppression with only minimal optimisation. This gives us reason to be optimistic about the levels of suppression that can be achieved in future work.
@@
The large magnetic pulses introduced in the PANCo scheme are also a significant potential noise source. We expect that variations in pulse strength will cause variations in the signatures, and thus cross-talk between magnetic and rotation signals similar to variation in bias field. In addition, these large pulses could induce eddy currents in nearby conductors such as magnetic shields, which may in turn cause magnetic noise that persists during the free precession period, and indeed the induced response in shields can approach 1 ms. This could vary due to changes in pulse size or external effects such as temperature changes, and so cause real and unwanted variations in magnetic fields. At present neither of these effects seem dominant in our experiment: we conclude this from a simple test with 10$\times$ weaker pulses (area of $\pi/20$, obtained by turning down the amplifier). This gave no corresponding reduction in low frequency magnetic noise floor, and only a degradation in rotational noise floor due to degraded magnetic suppression. We note that the coils used at present are deliberately small compared to the size of the shield (7.5 mm radius compared to $\approx 10$ cm distance to shields), resulting in a relatively small $\sim 100$ nT field at the shields during the pulses. We expect that future works will need careful design of coils to balance homogeneity on the sample while reducing eddy currents in the shields, in addition to careful tuning and control of the magnetic pulses. It is perhaps salient to note that the application of phase-controlling pulses are used in quantum information processing and NMR for dynamic decoupling [28, 35, 24]. While the purpose of the pulses in our system is different, there is a wealth of experimental and theoretical work that shows pulse sequences can be used in quantum limited systems with great success.
@@
Before probing the sensitivity limits outlined above, modifications will be required to our proof-of-principle experiment to address sources of noise and drift. Immediate improvements will come from replacing the current source used to drive the transverse coils. Following this, improvements to the stability of our optical beam paths to reduce pointing noise, pump beam inhomogeneity, optimising the gas cell mixtures and addressing the impact of field gradients will form a part of future experimental work.